OpenStudy (anonymous):

2s/s^2-25 + s/s-5 2s/s^2-25 + s/s-5 @Mathematics

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

(2/s)-9

OpenStudy (anonymous):

-29****

OpenStudy (anonymous):

yes add then simplify. ive worked it twice and gotten 2 different answeres

OpenStudy (anonymous):

\[\frac{2s}{s^2-25}+\frac{s}{s-5}=\frac{2s}{s^2-25}+\frac{s}{s-5}\times \frac{s+5}{s+5}\] \[=\frac{2s+s(s+5)}{s^2-25}\] then algebra in th numerator

OpenStudy (anonymous):

ok ive gotten that part, is that my answer? I dont know if i have to simplify more than that